# PLANesT-3D: A new annotated dataset for segmentation of 3D plant point clouds

### v1

@@ 4 Discussion
The PLANesT-3D dataset is composed of 3D point clouds of plants reconstructed through structure from motion and multi-view stereo. Capturing 3D geometry through SfM has many advantages such as low cost setup requirements, high resolution, and provision of color information [22]. It can be used both in controlled environments and in the field without extensive training for acquisition and elaborate protocols. A public dataset obtained through this modality as opposed to laser scanning is expected to be instrumental in assessing the robustness of 3D plant processing and phenotyping tools.
As can be observed in Table 3, the size of each model in terms of number of points is high, ranging from 600K to 5.5M points. One question is whether such high resolution is necessary for certain tasks related to plant analysis and management. The number of points can be reduced systematically through downsampling to measure the effect of resolution on the performance of particular 3D shape processing tools [48].
@@
We also used quite large numbers of images of the plants for reconstruction (2). This amount of image capture might not be feasible in cases where manual data acquisition is performed. However, advances in robotics enable fast acquisition with autonomous crop surveying robots equipped with robotic arms and multiple cameras [8, 49, 50], as well as techniques for automatic positioning of the robot arms for optimal data acquisition [51]. Nevertheless, limiting the number of images for reconstruction without a performance degradation in particular tasks would substantially reduce the computational demand.
Another important aspect of the new dataset is the availability of color information. Although not demonstrated in this work, use of color information together with 3D geometry can boost the performance of a segmentation procedure [52]. The PLANesT-3D dataset will be a valuable resource for measuring the contribution of color information for automated plant analysis and phenotyping tasks.
@@
Semantic parts annotated in the PLANesT-3D dataset are limited to "leaf" and "stem" classes; however, a more fine-grained annotation is possible, for example, in terms of "leaf blades", "petioles", "main stem", "nodes", and "internodes" [53]. Leaflets, together with petioles, can be grouped into their corresponding individual leaves. The architecture of each plant can be modeled with a graph representation. We envision undertaking this level of annotation as a future work.
@@
Another possible future work is a study on the transferability of the neural network models learned from a plant species into another species. In the current work, we trained a separate network for each of the three plant species. The extent of generalization ability of network models among plant species is of special interest; since manual annotation of plants for training is highly labor-intensive. PLANesT-3D, with a collection of three different species, allows application and evaluation of algorithms such as small sample learning and, transfer learning.
@@ 5 Conclusions
We introduced the PLANesT-3D dataset, which is composed of annotated color point clouds of 34 plants belonging to three different species. We described the acquisition, reconstruction, and labeling processes involved in dataset construction. As a use case of the dataset, we reported semantic segmentation results yielded by PointNet++, RoseSegNet, and a novel method abbreviated as SP-LSCnet. RoseSegNet achieved highest segmentation performance, reaching 95.5 %, 92.0 %, and 95.1 % MIoU, for pepper, rose, and ribes plants, respectively. SP-LSCnet also produced comparable results as 95.0 %, 89.8 %, and 94.5 % MIoU on the three plant species. We illustrated sample cases where the three methods failed to produce correct results. The PLANesT-3D dataset will be instrumental for developing new methods that effectively address these challenges.

### v2

@@ 4 Discussion
The main outcome of this study is the new PLANesT-3D dataset, which is composed of 3D point clouds of plants reconstructed through structure from motion and multi-view stereo. Capturing 3D geometry through SfM has many advantages such as low cost setup requirements, high resolution, and provision of color information [harandi2023make]. It can be used both in controlled environments and in the field without extensive training for acquisition and elaborate protocols. A public dataset obtained through this modality as opposed to laser scanning is expected to be instrumental in assessing the robustness of 3D plant processing and phenotyping tools.
PLANesT-3D is small in terms of the number of point clouds (34 point clouds) as compared to Pheno4D (126 point clouds) and Soybean-MVS (102 point clouds). It also lacks temporal data, i.e., all plants in PLANesT-3D were captured once and were not tracked during their development as opposed to the plants in Pheno4D (14 plants were tracked) and Soybean-MVS (5 plants were tracked). However, each of the point clouds of PLANesT-3D corresponds to a distinct plant with diverse overall shapes, architectures, and organ shapes and texture.
Obtaining 3D geometry through SfM has its limitations. One is the necessity of acquiring sharply-focused images from multiple locations. Each point on the object surface should be observed from at least two camera viewpoints for complete coverage. The manual acquisition process followed in this study may fall below the desired speed for high-throughput phenotyping. Furthermore, SfM is computationally intense compared to 3D geometry capture from Laser scanners or depth cameras, especially for generating dense, high-quality point clouds from large sets of images.
@@
We used quite large numbers of images of the plants for reconstruction (2). This amount of image capture might not be feasible in cases where manual data acquisition is performed. However, advances in robotics enable fast acquisition with autonomous crop surveying robots equipped with robotic arms and multiple cameras [atefi2021, iqbal2020development, smitt2021pathobot], as well as techniques for automatic positioning of the robot arms for optimal data acquisition [wu2019plant]. Nevertheless, limiting the number of images for reconstruction without a performance degradation in particular tasks would substantially reduce the computational demand.
We provided the 2D color images that were used to reconstruct the plant models as part of the PLANesT-3D dataset. As part of a future scenario, these images can be used to test new 3D reconstruction techniques that use 2D photographs as input, in terms of quality and computational efficiency.
As can be observed in Table 3, the size of each model in terms of number of points is high, ranging from 600K to 5.5M points. One question is whether such high resolution is necessary for certain tasks related to plant analysis and management. For all the semantic segmentation techniques used in this study, the number of points were reduced through downsampling as a preprocessing step. As future work, the downsampling rate can be varied systematically to measure the effect of resolution on the performance of particular 3D shape processing tools [Chaudhury2021transferring]. Another possibility is to search for optimal subsets of the images we provided for 3D reconstruction to reduce point density without loss of significant geometric information due to self-occlusion.
A weakness of this study is the lack of fully exploring the potential contribution of the color despite the availability of color information in the PLANesT-3D point clouds. The use of color information together with 3D geometry can boost the performance of a segmentation procedure [Boogaard2021boosting]. The potential boost depends on the plant species. For some plants, leaves and branches are indistinguishable in terms of color, for others color is one of the dominant attributes. An analysis on the contribution of color information deserves an extensive study. In this future analysis, the use of color correction schemes, conversion to various color spaces, and effects of illumination and plant species should be thoroughly explored.
The 3D color point clouds in PLANesT-3D, together with the corresponding color images, will form a valuable dataset, for which the acquisition was performed under uncontrolled illumination conditions. Also, color bleeding from the background in areas where there are abrupt changes in depth, i.e. boundaries of the organs, is present in the point clouds of the PLANesT-3D dataset. This issue needs to be resolved by methods aiming to integrate color information into the segmentation process.
We limited the study to the segmentation of isolated plants into their semantic parts. The evaluated techniques were set to operate under the condition that the plant point cloud was separated from the background through a prior step, and the correct scale of the plant was recovered. Segmentation of plants in cluttered environments, where object occlusions are also present, is beyond the scope of this study. However, we provide, as part of PLANesT-3D, the raw point clouds produced by Metashape Professional; i.e. the original point cloud prior to the application of our semi-automatic plant cropping and scaling procedure. These point clouds can be used to test techniques specifically designed to automatically separate plant points from the cluttered background, to recover correct scale, and to provide pose normalization.
@@
Semantic parts annotated in the PLANesT-3D dataset are limited to "leaf" and "stem" classes; which we believe is another limitaion of this study. A more fine-grained annotation is possible, for example, in terms of "leaf blades", "petioles", "main stem", "nodes", and "internodes" [Boogaard2022]. Leaflets, together with petioles, can be grouped into their corresponding individual leaves. The architecture of each plant can be modeled with a graph representation. We plan to undertake this level of annotation as a future work.
Another limitation of this study is the lack of application and comparison of panoptic segmentation algorithms, despite the availability of instance labels of the leaves. SP-LSCnet has the potential of being modified as an instance and panoptic segmentation method since it already performs a clustering of 3D points in its first stage. Once semantic labels are predicted in the second stage, these clusters can go through a proximity and homogeneity analysis for being merged into individual leaf instances. In a future study, the performance of this version of LSCnet can be compared with state-of-the-art panoptic segmentation networks extensively on PLANesT-3D datset and other datasets providing instance labels such as Pheno4D and Soybean-MVS.
We can list three more outcomes of this study as follows: 1) The semantic segmentation results on the new dataset obtained through the widely-used PointNet++ will serve as baselines for future studies. 2) RoseSegNet, without the necessity of tuning its hyperparameters, was able to surpass the segmentation performance of PointNet++. 3) SP-LSCnet, which is an alternative to end-to-end networks, provided competitive segmentation results. The MIoU results for each of the plant species were close with the three methods: around 95% for Pepper and Ribes, and around 90% for Rose.
The accuracy of semantic segmentation is particularly important, since semantic segmentation is either a prior step ([GUO2023]) or a parallel branch or module ([Vu2024, Li2022]) to instance or panoptic segmentation procedures. There is a significant imbalance between the number of points corresponding to the leaf and stem points due to the nature of the plants. The IoU for leaf points are high while, the IoU for stem is below 93% with all the methods. Correctly identifying the branches is particularly important for both panoptic segmentation and modeling the architecture of the plant. There is room for improvement in terms of both the precision and recall of the stem points, especially for the rose plants.
@@
Furthermore, with networks trained with other datasets of the same species, or with models from different species, one expects that the semantic segmentation results will drop below the baseline. The semantic segmentation results provided in this study, where a separate network was trained for each of the three plant species, can represent the achievable baselines for future studies on the transferability of the neural network models. The extent of generalization ability of network models among plant species is of special interest; since manual annotation of plants for training is highly labor-intensive.
In this regard, PLANesT-3D is important to the community in terms of species diversity, with a collection of three different species. The low-cost acquisition modality also contributes greatly to the data variation that will allow to evaluate algorithms for reproducibility, generalization ability, and suitability for transfer learning, and small sample learning.
This study also involves the introduction of the new technique, SP-LSCnet, which is composed of two modules. The first module, through t-SNE, extracts clusters of the 3D point clouds while visualizing the clusters in a single 2D layout. This visualization provides interpretability to the pipeline, as opposed to end-to-end deep learning architectures that operate as black-boxes. The second module can be replaced by any object classification network that processes 2D or 3D point clouds; allowing the method to be updated by more efficient methods in the future. The pipeline is also suitable for an interface where the user can visualize misclassified clusters in 2D and correct them if necessary.
Although PointNet++, RoseSegNet and SP-LSCnet seem to be performing similarly in terms of quantitative measures, the errors they produce have different natures (Figs. 13, 14, and 15). PointNet++ and RoseSegNet, in many cases, misidentify borders of leaves as stems. With SP-LSCnet this type of error is rare, as explained in Section 3. Both methods tend to misclassify thin petioles between close leaflets of the roses; however, while PointNet++ and RoseSegNet’s errors are scattered individual points, SP-LSCnet misclassifications appear as smooth regions in between leaflets. Other sources of error for SP-LSCnet are the misidentification of small clusters produced by t-SNE at branch junctions and regions of low point density. Including context information using multi-scale clustering with varying perplexity values, which are then organized in a graph structure can help reduce such errors of the SP-LSCnet framework.
One weakness of SP-LSCnet is the reliance on the semantic homogeneity of superpoints, which is not always guaranteed, especially at leaf-stem boundaries. Another, as mentioned above, is the loss of context, since each superpoint is processed individually. Our future work involves remedying these weaknesses by processing multiple t-SNE maps in a hierarchical manner, organizing superpoints in graph structures, and using graph networks for classifying each node into its semantic class.
The effectiveness of SP-LSCnet, should also be tested on other 3D plant datasets mentioned in Table 1. Furthermore, cross-validation tests should be performed, where the network is trained on one dataset (or species) and tested on another. In order to keep this current study focused, an extensive cross-validation study of the current SP-LSCnet and its improved versions is reserved to be part of a future work.
@@ 5 Conclusions
We introduced the PLANesT-3D dataset, which is composed of annotated color point clouds of 34 plants belonging to three different species. We described the acquisition, reconstruction, and labeling processes involved in dataset construction. PLANesT-3D will be a valuable benchmark for assessing the generalization ability and reproducibility of current and future 3D plant phenotyping tools on diverse data. PLANesT-3D includes new instances from three different species. Point clouds of two of these species are provided for the first time. The instances posses a large within-species variety in terms of plant size, architectural complexity, and the number, density, geometry and texture of organs. Also, the new dataset is constructed with the low-cost SvM/MVS technique, introducing challenges such as varying point density, missing data, reconstruction noise, and color bleeding.
As a use case of the dataset, we reported semantic segmentation results yielded by PointNet++, RoseSegNet, and a novel method abbreviated as SP-LSCnet. RoseSegNet achieved highest segmentation performance, reaching 95.5 %, 92.0 %, and 95.1 % MIoU, for pepper, rose, and ribes plants, respectively. SP-LSCnet also produced comparable results as 95.0 %, 89.8 %, and 94.5 % MIoU on the three plant species. The main advantage of SP-LSCnet is its two-stage structure, where the first stage is an unsupervised scheme that provides 2D visualization and the second stage can be replaced by any classification network. The 2D visualization greatly increases the interpretability of the method in comparison to state-of-the-art black-box techniques.
We illustrated visual errors of the three methods in their semantic segmentation outputs. The PLANesT-3D dataset will be instrumental for developing new semantic segmentation methods that effectively address the challenges illustrated by these errors. As a dataset containing plant point clouds of diverse structures with regions of varying quality, it will help assessing tools for 3D plant analysis tasks, such as panoptic segmentation, leaf counting, leaf shape characterization, skeletonization of the branch structure, decoding of the architecture, which are essential for automating plant research and agricultural operations.
@@ Acknowledgements:
Appendix A LSCnet
In this section, we provide details on the Leaf-Stem-Classification network (LSCnet), which classifies individual superpoints (Fig. 10). The network is formed through integrating the Center Shift Module (CSM) and the Radius Update Module (RUM) [turgut2023local] to the PointNet++ architecture for object classification (Fig. 16). The modules update adaptively the centers and radii of the spherical regions on which PointNet++ encodes local features.
Figure 16: The integration of CSM and RUM to PointNet++ framework [turgut2023local]: The classification network. $r^{L}$ for $L=1$ (first layer) is set to be 0.2. $r^{L}$ for $L=2$ (second layer) is 0.4 (a). Adaptive local region inference with CSM and RUM (b).
The classification network architecture of PointNet++ [qi2017pointnet2], shown in Fig. 16(a), includes three set abstraction (SA) layers. The input point cloud with $N=1024$ points is reduced to $N_{1}=512$, and $N_{2}=128$ representative points during SA layer-1 and SA layer-2, respectively. At each SA layer $L$, the representative points $c_{j}^{L}$ are selected using farthest point sampling (FPS) algorithm. $N_{L}$ spherical local regions with radius $r^{L}$ is defined around the center points $c_{j}^{L}$. Within each region, $K$ points are randomly selected and the features of $K$ points are mapped to higher dimensions with MLPs. The abstracted features of each local region are calculated by taking the maximum among the feature channels of $K$ grouped points.
The adaptive local inference modules, CSM and RUM can be integrated to either or both of the first two layers as shown in Fig. 16(b). In other words, CSM and RUM modules can be turned "ON" or "OFF" at particular layers of the network. If CSM module is set as "ON", the representative points are updated. If it is "OFF", representative points remain as the ones computed by the FPS algorithm. Likewise, if the RUM module is "ON", the radius around each representative point is altered, otherwise it is set to the constant value specific to each layer. In this way, the user can control at which layers the receptive fields are altered adaptively using attention mechanisms.
If CSM is set as "ON" at a particular layer, for each representative point $c_{j}^{L}$, CSM calculates the shift amount $\Delta c_{j}^{L}$, through attention-based subnetworks. The representative point is updated as:
$\hat{c}_{j}^{L}=c_{j}^{L}+\Delta c_{j}^{L}\;.$ (4)
$\hat{r}_{j}^{L}=r^{L}+\Delta r_{j}^{L}\;.$ (5)
Various alternatives of CSM and RUM were suggested by [turgut2023local]. As a result of experimental trials, the CSM-II (sub) and RUM-II (cum) versions were selected for the PLANesT-3D dataset. Please, refer to [turgut2023local] for detailed descriptions of the modules CSM-II (sub) and RUM-II (cum).